OpenStudy (anonymous):

While trying out for the position of pitcher on your high school baseball team, you throw a fastball at 84.9 mi/h toward home plate, which is 18.4 m away. How far does the ball drop due to effects of gravity by the time it reaches home plate? (Ignore any effects due to air resistance and assume you throw the ball horizontally.) ==>_______m

OpenStudy (anonymous):

@Jemurray3

OpenStudy (anonymous):

The first thing you need to know is how long the ball was in the air. To do that, you should divide the distance it travels by its horizontal velocity. However, the velocity and position are in incompatible units -- first convert the velocity to m/s.

OpenStudy (anonymous):

so the conversion is 37.95 m/s

OpenStudy (anonymous):

To find the time it takes to go from the pitcher's mound to the plate, divide the distance travelled by the velocity.

OpenStudy (anonymous):

so 18.4/37.95?

OpenStudy (anonymous):

@Jemurray3

OpenStudy (anonymous):

Yes. Now that you have the time, you know that the distance it falls in the y direction is given by the equation \[ \Delta y = -\frac{1}{2} gt^2 \] where g is the gravitational constant 9.81 m/s^2.

OpenStudy (anonymous):

so it would equal 2.4?

OpenStudy (anonymous):

No. Check your work.

OpenStudy (anonymous):

i forgot the negative.... so -2.4?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

change in y= -1/2*9.81*0.49^2 = -1.18?

OpenStudy (anonymous):

@Jemurray3

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

that was apparently wrong though ;/

OpenStudy (anonymous):

if it's asking how far the ball drops, perhaps it would like you to disregard the minus sign.

OpenStudy (anonymous):

so a positive 1.18?

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

still wrong...

OpenStudy (anonymous):

Using the proper number of significant figures gets me 1.15 meters.

OpenStudy (anonymous):

that was correct. did I calculate wrong?

OpenStudy (anonymous):

don't round your answers in the middle of the calculation.

OpenStudy (anonymous):

oh i seee. okay.. Thank you!!!(: